business finance

Big Steve's, makers of swizzle sticks, is considering the purchase of a new plastic stamping machine.  This investment requires an initial outlay of

$

and will generate net cash inflows of

$

per year for

years.

a.What is the project's NPV using a discount rate of

percent?

Should the project be accepted?  Why or why not?

b.What is the project's NPV using a discount rate of

percent?  Should the project be accepted?  Why or why not?

c.What is this project's internal rate of return?  Should the project be accepted?  Why or why not?

a.If the discount rate is

percent, then the project's NPV is
Answer #1

CALCULATION OF NET PRESENT VALUE =

PRESENT VALUE OF CASH INFLOW - PRESENT VALUE OF CASH OUTFLOW

PRESENT VALUE OF CASH INFLOW = 4.34*20000 =86860

PRESENT VALUE OF CASH OUTFLOW = 105000

NET PRESENT VALUE = 105000-86860

= (18140)


answered by: ANURANJAN SARSAM
Answer #2

SOLUTION :


a.


Initial cost = 90000 ($)

Discount rate , r = 9% = 0.09

=> (1 + r) = 1.09

Project generates $17000 per year for 9 years.


So,


NPV 

= - Initial cost + PV of future cash flows

= - 90000 + 17000(1.09^9 - 1)/(0.09*1.09^9)

= 11919.20 ($) (ANSWER).

Project is acceptable as NPV is positive. (ANSWER).


b.


If discount rate, r = 14% = 0.14

=> (1+r) = 1.14


So,


NPV 

= - Initial cost + PV of future cash flows

= - 90000 + 17000(1.14^9 - 1)/(0.14*1.14^9)

= - 5911.68 ($) (ANSWER).

Project is not acceptable as NPV is negative. (ANSWER).


c.


At IRR, NPV = 0 .

Let IRR be r (in decimals).

So, NPV = 0 = 17000((1+r)^9 - 1)/(r(1+r)^9) 

By trial and error, r = IRR = 0.1216938 = 0.1217 approx. 

So, IRR is = 12.17% (ANSWER).


If discount rate is 9%, IRR is higher and project can be accepted.

If discount rate is 14%, IRR is lower and project should not be accepted. (ANSWER).


answered by: Tulsiram Garg